PHOTOCHEMISTRY OF NITROMETHANE ADSORBED ON ALKALI HALIDE FILMS

TONI GREEN, \underline{ANGELA JARAMILLO}, GERALD WAYTON, and C.A. BAUMANN, Department of Chemistry, University of Scranton, Scranton, PA 18510.

Nitromethane was adsorbed at low temperature onto sublimated alkali halide films. Irradiation of the adsorbate at wavelengths below 250 nm produced a variety of photoproduct species. We have identified a number of the primary photoproducts (CH_{3}ONO, H_{2}CO, HNO) as well as species resulting from secondary photolysis (CO, N_{2}O). The quantum efficiencies for production of these species were determined as a function of irradiation wavelength and surface temperature.